OpenStudy (anonymous):

After a certain drug is injected into a patient, the concentration c of the drug in the bloodstream is monitored. At time t ≥ 0 (in minutes since the injection), the concentration (in mg/L) is given by c(t) =(40t)/(t^2 + 2) What eventually happens to the concentration of drug in the bloodstream?

OpenStudy (anonymous):

it deceases to 0

OpenStudy (anonymous):

since degree of the denominator (2) is bigger than degree of the numerator (1) \[\lim_{t\rightarrow \infty} c(t)=0\]

OpenStudy (anonymous):

would that be the same as a \[-\infty\]

OpenStudy (anonymous):

ooooooooooooooooh no

OpenStudy (anonymous):

0 is 0 not \[-\infty\]

OpenStudy (anonymous):

ok just making sure because i knew your answer made sense i had just never had a problem like that with an answer 0

OpenStudy (anonymous):

doesn't make any sense in this case anyway. you cannot have a negative amount of a drug in your blood, let alone - infinity amount!

OpenStudy (anonymous):

yeah lol i just usually had the option between \[-\infty or \infty\]

OpenStudy (anonymous):

so i just wanted to make sure